Grade 12th passMechanics

Four forces 180N,100N,60N, and 50N are acting along sides AB,BC,CD, and DA of a square ABCD of side 2m.Their resultant force is 130N.Calculate position of resultant w.r.t A

Rohit
6 Years agoGrade 12th pass

4 Answers

Venkat
6 Years ago
Take moments about point ABC x 2m +CD x 2 m = Reslutant x yy is the unknown distance from A.100 N x 2m + 60 N x 2m = 130 N x y m y = 320/130 = 2.46 m
Rugved Patil
5 Years ago
Four forces 180N,100N,60N, and 50N are acting along sides AB,BC,CD, and DA of a square ABCD of side 2m.Their resultant force is 130N.Calculate position of resultant w.r.t A
 
Patel hetal
5 Years ago
Four forces of magnitude 50 N, 100 N, 80 N and 140 N are acting along the sides
of a squire AB, BC, CD, and DA respectively. The resultant force is given by.
Aditya
4 Years ago
Four forces 180N, 100N, 60N, and 50N are acting along sides AB, BC, CD, and DA of a square ABCD of side 2m. Their resultant force is 130N. Calculate position of resultant w.r.t A *